The mass of magnesium hydroxide, Mg(OH)2, in an indigestion tablet was determined as

follows:

Step 1 The tablet was crushed and dissolved in exactly 40.00 cm3

of 0.5M


Hydrochloric acid (an excess).

Mg(OH)2(s) + 2HCl(aq)  MgCl2(aq) + 2H2O(l)

Step 2 The amount of hydrochloric acid remaining was measured by titration with

0.250 mol dm–3


sodium hydroxide solution.


22.80 cm3


of this sodium hydroxide solution was required.


Calculate the mass of magnesium hydroxide in the tablet.


1
Expert's answer
2021-03-11T08:42:14-0500

total mmol of HCl = 40×0.5 = 20 mmol

mmol of HCl remaining = 0.250×22.80 = 5.7

mmol

moles of HCl reacted with Mg(OH)2 =

20 - 5.7 = 14.3 mmol

mmol of Mg(OH)2 required = 14.3/2 = 7.15 mmol

mass of Mg(OH)2 = (7.15/100)×58.3 = 0.4168g